You illuminate a slit with a width of 0.0605 mm with a light of wavelength 719 nm and observe the resulting diffraction pattern on a screen that is situated 2.33 m from the slit. What is the width, in centimeters, of the pattern's central maximum?
Answer #1

Solution)

We know,

The angle to the first minimum

sinθ1min=mλ/D =1*719*10^-9/0.0605*10^-3=0.0118

Now,

Y=Lsinθ =2.33*0.0188=0.0276 m

Now, width of the central maximum is

W=2Y=2*0.0276=0.0553

W=5.53 cm

============